OpenStudy (anonymous):

Heather's family bought tickets at a local carnival. Her brother took1/5 the number of tickets that her cousin took, which was half the number Heather took. If her brother took 10 tickets, how many tickets did Heather take? 10 25 50 100

OpenStudy (anonymous):

B = (1/5)C C = (1/2)H So, B = (1/5)(1/2)H B = (1/10)H -> (10)B = H (10)(10) = H = 100

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

All good now, @Yasminelove1 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @Yasminelove1 Yes! 100

OpenStudy (anonymous):

no prob thank you u helped me out a whole lot

OpenStudy (anonymous):

You're nice to work with!

OpenStudy (anonymous):

can u help me with this Kyle drives 2 4/5 hours to visit his grandmother when he takes the new expressway, which has made his trip 1/3 as long as it used to be. How long did his trip take before the expressway was built? 12/15 hours 3 2/15 hours 8 2/5 hours 2 7/15 h ours

OpenStudy (anonymous):

The new trip is 2 and 4/5 hours which is 2 hours and 48 minutes. Since this 1/3 of the original time, the original time is: (2)(3) hours and (3)(48) minutes = 6 hours and 144 minutes or 8 hours and 24 minutes which is: 8 and 2/5 hours

OpenStudy (anonymous):

Got it now, @Yasminelove1

OpenStudy (anonymous):

yes a little

OpenStudy (anonymous):

Renee took a long flight to visit her family that lives out of the country. She slept for the first half of the trip and watched movies for one-fourth of the remaining flight. If the total trip took ten and one-half hours, how long did Renee watch movies? 1 hour one and five-sixteenths hours 2 hours one and one-fourth hours

OpenStudy (anonymous):

i really dont understand this one at all

OpenStudy (anonymous):

ok, last one and then I have to go after this one. Have to look after a senior. btw, you did get the answer to the second one didn't you? I posted it.

OpenStudy (anonymous):

She watched movies for 1/4 of the remaining 1/2 flight or: 1/8 of the flight (10.5)(1/8) = 21/16 = 1 and 5/16

OpenStudy (anonymous):

All done! @Yasminelove1

OpenStudy (anonymous):

thank you so much yah the best

OpenStudy (anonymous):

Okay! You now have answers to all 3 questions! Good luck to you!

OpenStudy (anonymous):

uw!

OpenStudy (anonymous):

i suck in math ur a life saver

OpenStudy (anonymous):

But I'm a terrible artist!

OpenStudy (anonymous):

oh wow i am to dont feel bad at least u good in math